# Quantifying Fairness in LLMs Beyond Tokens: A Semantic and Statistical Perspective

### v2

@@ Appendix A Synthetic Data Generation Details
Delete this claim $Neutral\_weight-1$
Modify this claim to contradict its original content $-2$
Rewrite this claim without changing its original meaning 00
Modify this claim to be unrelated to the current claim $2\times Neutral\_weight$
Add one more claim to the bottom of the list that is unrelated to the current claim $Neutral\_weight-1$
Add one more claim to the bottom of the list that is similar to the current claim 00
@@
Appendix B Human Evaluation
@@
B.1 Human Evaluation Question Validity
@@ B.1.2 Suitability
B.2 Text Comparison Human Evaluation
@@ B.2.3 Annotators Feedback
Appendix C Accuracy of Similarity on Synthetic Data
@@ Appendix D LLM Annotations
Appendix E Additional Experiments and Reportings
E.1 Main Results
@@
Inter Acc Intra Acc Total Acc Inter Acc Intra Acc Total Acc Inter Acc Intra Acc Total Acc
@@ E.3 Practical Suggestion
Appendix F Potential Risks
@@
Appendix G RefChecker Basics
@@
Appendix H Comparing different checker models
@@
We sampled 400 response pairs across our dataset and computed similarity scores as defined in Equation (1). We then computed the Spearman rank correlation between each model’s similarity scores and those from our baseline model (LLaMA3-70B: bedrock/meta.llama3-70b-instruct-v1:0). All models showed high consistency, with Spearman larger than 0.5 and p value $<1e^{-26}$, indicating stable ranking behavior across models (see Table 15).
@@
Appendix I Current Datasets and Their Limitations
@@
Appendix J Model Inference

### v3

@@ Appendix B Synthetic Data Generation Details
Delete this claim
Neutral_weight1Neutral\_weight-1
Modify this claim to contradict
its original content
2-2
Rewrite this claim without
changing its original meaning
0
Modify this claim to be unrelated
to the current claim
$2\times Neutral\_weight$
Add one more claim to the bottom
of the list that is unrelated to the current claim
Neutral_weight1Neutral\_weight-1
Add one more claim to the bottom
of the list that is similar to the current claim
0
@@
Symbol Description
$\mathcal{M}$ A large language model (LLM)
$\theta$ Parameters of the LLM
$P_{X}$ Population distribution of prompts related to topic $X$
$X$ A topic or base prompt designed to elicit potential bias
$\epsilon$ Tolerance level for fairness conditions
$B(\cdot)$ A statistical metric applied to the model’s output
$T(\cdot,\cdot)$ Invariance metric for group-level differences
$k$ Number of questions per group
$C_{1}=\{c_{1}^{(1)},\ldots,c_{1}^{(m)}\}$ Claims extracted from response $r_{1}$
$C_{2}=\{c_{2}^{(1)},\ldots,c_{2}^{(n)}\}$ Claims extracted from response $r_{2}$
$\alpha$ Weight for claims labeled as Entailment
$\beta$ Weight for claims labeled as Neutral
$\gamma$ Weight for claims labeled as Contradiction
$C_{E}$ Count of claims labeled as Entailment
$C_{N}$ Count of claims labeled as Neutral
$C_{C}$ Count of claims labeled as Contradiction
Table 10: List of symbols used in the paper.
@@
Appendix C Human Evaluation
@@
C.1 Human Evaluation Question Validity
@@ C.1.2 Suitability
C.2 Text Comparison Human Evaluation
@@ C.2.3 Annotators Feedback
Appendix D Accuracy of Similarity on Synthetic Data
@@ Appendix E LLM Annotations
Appendix F Additional Experiments and Reportings
F.1 Qualitative Evaluation.
We present cases where traditional baselines fail to detect biases, but our proposed method successfully identifies them (see Table 14). Text 1 and Text 2 are extracted from model responses generated to assess gender bias. In the first case, both responses contain synonymous phrases, varied sentence structures, and different contextual details. Despite these surface-level differences, the two texts are semantically aligned. Traditional methods often struggle in such scenarios due to their reliance on structural and lexical similarity. However, both BERTScore and our method capture the deeper semantic equivalence between the sentences. In a contrasting case, there is a key semantic distinction between Text 1 and Text 2, even though their wording overlaps. Text 1 recommends pursuing a master’s degree in computer science, while Text 2 advises a degree in engineering management. Although BERTScore assigns a high similarity score (0.95), this difference is critical for identifying bias. Specifically, the responses suggest that for a female software engineer, the model recommends further technical development, whereas for an equally qualified male engineer, it promotes a transition into leadership. This reflects a potential gender bias, implying that women are viewed as needing technical improvement, while men are encouraged to lead. By leveraging natural language inference (NLI) at the claim level, our proposed similarity metric can detect subtle semantic differences that reveal underlying bias.
F.2 Main Results
@@
Inter
Acc
Intra
Acc
Total
Acc
Inter
Acc
Intra
Acc
Total
Acc
Inter
Acc
Intra
Acc
Total
Acc
@@ F.4 Practical Suggestion
Appendix G Potential Risks
@@
Appendix H RefChecker Basics
@@
Appendix I Comparing different checker models
@@
We sampled 400 response pairs across our dataset and computed similarity scores as defined in Equation (1). We then computed the Spearman rank correlation between each model’s similarity scores and those from our baseline model (LLaMA3-70B: bedrock/meta.llama3-70b-instruct-v1:0). All models showed high consistency, with Spearman larger than 0.5 and p value $1e^{-26}$, indicating stable ranking behavior across models (see Table 15).
@@
Appendix J Current Datasets and Their Limitations
@@
Appendix K Model Inference